Given the following data, calculate ∆G for this chemical reaction, CH4 + 2O2 → CO2 + 2H2O, if H = 750 J , S = 25.75 J/K, and T = 315 K.


1
Expert's answer
2021-06-25T05:41:27-0400

ΔG=ΔH−TΔS.

= 750 -(25×315)

= 750-(7875)

= -7125J K‑1 mol-1